# A High-Speed Hardware Algorithm for Modulus Operation and its Application in Prime Number Calculation

### v1

@@ 4.3 Prime Number Calculation
Form the Figure 8 we can observe that the Python implementation shows significantly longer computation times compared to both the hardware and C implementations, while C implementation performs better than hardware implementation. Notably, the hardware algorithm on Zybo FPGA is currently operating at 125 MHz, but it has potential to operate at a maximum frequency of 295.5 MHz. If the hardware were run at this higher frequency, the performance could potentially surpass that of the C implementation on the high-performance PC, particularly for lager integer values.
@@ 4.4 Comparison
Our implementation of the modular exponentiation algorithm shows significant improvements over these previous works. For 2048-bit operand sizes, our design utilized only 45% of the available LUTs (17,600 total) on a Zybo FPGA board. Compared to Butler et al. [15], our implementation uses substantially fewer FPGA resources while achieving a higher operating frequency of 295.5 MHz. This represents a notable advancement in both resource efficiency and performance for modular arithmetic operations on FPGAs.
@@ 4.5 Discussion
The linear relationship between the bit length difference (BLD) and the cycle count, as demonstrated in Figure 7, highlights the efficiency of the proposed hardware algorithm for modulus operation. The algorithm scales predictably, maintaining efficiency even as operand sizes increase. This linear scalability is crucial for applications requiring high-speed computations with varying operand sizes.
Figure 7(d) provides additional validation of the algorithm’s performance by comparing experimental data with the predicted cycle count. The close alignment of the experimental data points with the predicted cycle count based on the equation $y=2x+2$ highlights the accuracy and robustness of the prediction model. The experimental results confirm that the hardware algorithm performs efficiently and predictably across different BLD values, maintaining low cycle counts even as the BLD increases.
@@
The results validate the efficacy of the proposed hardware algorithm for modulus operations and its application in prime number calculations. The linear scalability, efficient resource utilization, robust timing performance, and effective power management make the algorithm well-suited for a wide range of hardware applications. Future work could explore further optimization to increase the operating frequency and reduce power consumption, as well as adaptations for other computationally intensive tasks.
@@ 5 Conclusion
This study introduces a novel hardware algorithm for the modulus operation, optimized for FPGA implementation. The algorithm’s linear scalability and low overhead, demonstrated through extensive testing with operand sizes from 32-bit to 2048-bit, make it a robust solution for high-speed computations. The implementation on the Zybo FPGA platform confirms efficient resource utilization, robust timing performance, and effective power management, underscoring its suitability for both high-performance and resource-constrained platforms. Its application in prime number calculation further validates its practical use, showcasing substantial performance improvements over traditional software implementations. The findings suggest that this hardware algorithm can notably accelerate cryptographic protocols and other applications reliant on modular arithmetic, offering a promising direction for further research and development in hardware-based arithmetic operations.

### v2

@@ 4.3 Prime Number Calculation
Form the Figure 10 we can observe that the Python implementation shows significantly longer computation times compared to both the hardware and C implementations, while C implementation performs better than hardware implementation. Notably, the hardware implementation of the algorithm on Zybo FPGA is currently operating at 125 MHz, but it has potential to operate at a maximum frequency of 295.5 MHz. If the hardware were run at this higher frequency, the performance could potentially surpass that of the C implementation on the high-performance PC, particularly for lager integer values.
@@ 4.4 Comparison
Our implementation of the modular exponentiation algorithm shows significant improvements over these previous works. For 2048-bit operand sizes, our design utilized only 7,920(45%) of the available LUTs (17,600 total) on a Zybo FPGA board. Compared to Butler et al. [15], our implementation uses substantially fewer FPGA resources while achieving a higher operating frequency of 295.5 MHz. This represents a notable advancement in both resource efficiency and performance for modular arithmetic operations on FPGAs.
In addressing the limitations found in the existing literature, our proposed algorithm offers several key improvements. One of the primary challenges identified in prior works is the complexity and resource intensity of performing molar arithmetic, particularly when relying on operations such as multiplication and division. Our approach, by focusing on addition, subtraction, logical operations and bit shifts, avoid these costly operations and thereby reduces both the hardware complexity and the associated resource utilization. Furthermore, the scalability of our algorithm is demonstrated by its linear cycle count scaling with operand size, an aspect often neglected in existing implementations where non-linear growth is observed. This predictable scalability makes our design more suitable for cryptographic applications that require handling large operand sizes. Additionally, by optimizing the modulus operation itself, rather than focusing on composite operations such as modular multiplication or exponentiation, we provide a fundamental improvement that can be leveraged across modular arithmetic operations, further addressing the gaps in previous hardware designs.
@@ 4.5 Discussion
The linear relationship between the bit length difference (BLD) and the cycle count, as demonstrated in Figure 9, highlights the efficiency of the proposed algorithm for modulus operation. The algorithm scales predictably, maintaining efficiency even as operand sizes increase. This linear scalability is crucial for applications requiring high-speed computations with varying operand sizes.
Figure 9(d) provides additional validation of the algorithm’s performance by comparing experimental data with the predicted cycle count. The close alignment of the experimental data points with the predicted cycle count based on the equation $y=2x+2$ highlights the accuracy and robustness of the prediction model. The experimental results confirm that the algorithm performs efficiently and predictably across different BLD values, maintaining low cycle counts even as the BLD increases.
@@
The results validate the efficacy of the proposed algorithm for modulus operations and its application in prime number calculations. The linear scalability, efficient resource utilization, robust timing performance, and effective power management make the algorithm well-suited for a wide range of hardware applications. Future work could explore further optimization to increase the operating frequency and reduce power consumption, as well as adaptations for other computationally intensive tasks.
@@ 5 Conclusion
This study introduces a novel algorithm for the modulus operation, optimized for FPGA implementation. The algorithm’s linear scalability and low overhead, demonstrated through extensive testing with operand sizes from 32-bit to 2048-bit, make it a robust solution for high-speed computations. The implementation on the Zybo FPGA platform confirms efficient resource utilization, robust timing performance, and effective power management, underscoring its suitability for both high-performance and resource-constrained platforms. Its application in prime number calculation further validates its practical use, showcasing substantial performance improvements over traditional software implementations. The findings suggest that this algorithm can notably accelerate cryptographic protocols and other applications reliant on modular arithmetic, offering a promising direction for further research and development in hardware-based arithmetic operations.